OpenStudy (anonymous):

Find the equation of a parabola passing through (-4, -6) whose vertex is at (-2,-4) and whose axis is parallel to the x-axis. the answer is (y+4)^2 = -2(x+2).. how to get the P given the points, vertex? Help me please. Show it to me step by step. Thank you.

OpenStudy (anonymous):

vertex is \((-2,4)\) and the axis of symmetry is parallel to the \(x\) axis, so you know it looks like \[(y+4)^2=a(x+2)\] is that much clear ?

OpenStudy (anonymous):

oops i meant vertex is \((-2,-4)\)

OpenStudy (anonymous):

yes. and how about a? how to get a?

OpenStudy (anonymous):

for that you need to use the second point

OpenStudy (anonymous):

replace \(x\) by\(-4\) and \(y\) by \(-6\) and solve for \(a\)

OpenStudy (anonymous):

\[(y+4)^2=a(x+2)\] \[(-6+4)^2=a(-4+2)\]

OpenStudy (anonymous):

hooray! got it now!!

OpenStudy (anonymous):

super thank you :)))

OpenStudy (anonymous):

you get \[4=-2a\] and so \(a=-2\) yw!!